Come and Go Bank offers your firm a discount interest loan with an interest rate of 9 percent for up to $21 million, and in addition requires you to maintain a 2 percent compensating balance against the face amount borrowed.

What is the effective annual interest rate on this lending arrangement? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Effective annual rate= %
Answer #1

Interest Paid = 0.09(21) = $1.89 million

Amount Received = (21 - 1.89) - (21)(0.02)

Amount Received = $18.69 million

EAR = 1.89/18.69

EAR = 10.11%